Grade 11Physical Chemistry

If the no. of molecules of SO2 effusing through an orifice of unit area of cross section in unit time at 0 degree Celsius at 1 atm pressure is n , what is the no. of He molecules effusinf through similar conditions at 273 degree Celsius and 0.25 atm?

Soham
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
Dear Soham
 
r1/ r2 = (p1 / p2)(sqrt (T2 M2/ T1 M1))
 
n/ x = (1 /0.25) (1/sqrt(8))
 
x = n/ sqrt(2)
 
Hope it helps
 
Regards
Arun